# Recitation over Reasoning: How Cutting-Edge Language Models Can Fail on Elementary School-Level Reasoning Problems?

### v2

@@ B.1 Best-of-5 Results
Tab. 5 (for text-based problems) and Tab. 6 (for vision-based problems) shows the best-of-5 result of the experiments conducted in Sec. 4.1. The conclusion is very similar to those in Sec. 4.1, indicating that the problem is hard to fix with LLM alignment techniques such as reinforcement learning [52].
@@
B.2 Greedy Decoding Results
Tab. 7 (for text-based problems) and Tab. 8 (for vision-based problems) shows the average score of LLMs doing greedy-decoding (i.e. temperature=0) in the experiments conducted in Sec. 4.1. The conclusion is similar to those in Sec. 4.1.
@@ B.2 Greedy Decoding Results
B.3 More Results on Few-Shot In-Context Learning
Due to space limit, we only show the results of some most representative LLMs in Sec. 4.2; we show the results for the all tested LLMs in Tab. 9. We found that models with weaker base ability, such as Qwen2.5-7B-Instruct and Qwen2.5-14B-Instruct, are less benefited from few-shot ICL techniques.
@@ B.3 More Results on Few-Shot In-Context Learning
B.4 More Results on “No Solution” Problems
Due to space limit, we only show the results of some most representative LLMs in Sec. 4.3; we show the results for the all tested LLMs in Tab. 10. We find that while state-of-the-art models we reported in the main paper generally benefits from fixes such as adding “forced correct” prompt or 1-shot, many other relatively weaker models, including Qwen-2.5 small models, GPT-4o-mini-0718 and QwQ-32B-Preview, still struggles much on addressing “no solution” problems.
@@ B.4 More Results on “No Solution” Problems
Table 10: Results of all LLMs with the settings in Sec. B.4. While generally adding “forced correct” prompts and 1-shot helps the performance, most models still heavily struggle.
B.5 Failure Example on Original Problem with “Forced Correct” Prompt
In Sec. 4.1.1, the experiment results show that with “Forced Correct” prompt, the performance of LLMs on original problems generally decrease, and such issue is particularly severe with Gemini-2.0 Pro-0205. Here we show an example of the LLM getting confused with the prompt on original problems in the box below, showing that it is non-trivial to balance correctness and user experiences with the inputs that are not so rigorous.
# Problem{CJK} UTF8gkai 请回答下面的问题。题目保证没有错误,请严格按照题目字面所写的问题回答。 【问题】你正在图书馆枯坐,一位陌生美女主动过来和你搭讪,递给你两个相同的硬币并要求和你一起玩个数学游戏。美女提议:“让我们各自亮出硬币的一面,或花或字。如果我们都是花面,那么我给你 3 元,如果我们都是字面,我给你 1 元,剩下的情况你给我 2 元就可以了。” 那么,这个游戏公平吗? (Please answer the following problems. The problems are guaranteed to be correct; please strictly follow the literal meaning of the problem. [Question] You are sitting idly in the library when a beautiful stranger approaches you and strikes up a conversation. She hands you two identical coins and proposes a math game. She says: "Let’s each reveal one side of a coin—either heads or tails. If we both show heads, I’ll give you 3 yuan. If we both show tails, I’ll give you 1 yuan. In all other cases, you just give me 2 yuan." So, is this game fair?) # Gemini-2.0 Pro-0205 Response{CJK} UTF8gkai 这个游戏是否公平,取决于 “公平” 的定义是什么。题目没有问这个游戏是否公平,也没有定义什么是 “公平”,因此无法回答这个问题。 (Whether the game is “fair” depends on the definition of fair. The problem does not ask whether the game is fair, and did not define what is “fairness”; therefore, the problem cannot be answered.)
B.6 English Version of Fig. 2
The response for OpenAI-o1-1217 for our English translation of the text problems in Fig. 2 can be seen in https://chatgpt.com/share/67f45d45-a694-800d-a4c8-3c763e93400f. In later experiments, the response of GPT-4.5 points out that there is a possibility of ambiguity - “traveling in opposite directions” could also possibly mean traveling towards each other, as if the two cars are located in the west and east respectively, the car on the west will need to travel east and the car on the east will need to travel west (https://chatgpt.com/share/67f45e46-5eb0-800d-8f1c-5dec1f3c9f1d), which are indeed opposite directions. To further address such concerns, we test the result of literal translation from Chinese (“traveling back to back”), and the following version with much more explicit explanation:
Two cars start simultaneously from two cities that are 300 km apart, one on the east and the other on the west. The car on the east travels towards the east, and the car on the west travels towards the west. One car has a speed of 60 km/h, while the other has a speed of 70 km/h. How many hours will it take for them to meet?
Surprisingly, we found OpenAI-o1-1217 still fails on both the literal translation (https://chatgpt.com/share/67f4601a-149c-800d-8be1-d4ebe97c315b) and the explicit version above (https://chatgpt.com/share/67f444ed-8148-800d-810d-62c54ea1636a).

### v3

@@ C.4 Best-of-5 Results
Tab. 8 (for text-based problems) and Tab. 9 (for vision-based problems) shows the best-of-5 result of the experiments conducted in Sec. 4.1 and Sec. 4.2. The conclusion is very similar to those in Sec. 4.1 and Sec. 4.2, indicating that the problem is hard to fix with LLM alignment techniques such as reinforcement learning [53].
@@
C.5 Greedy Decoding Results
Tab. 10 (for text-based problems) and Tab. 11 (for vision-based problems) shows the average score of LLMs doing greedy-decoding (i.e. temperature=0) in the experiments conducted in Sec. 4.1 and Sec. 4.2. The conclusion is similar to those in Sec. 4.1 and Sec. 4.2.
@@ C.5 Greedy Decoding Results
C.6 More Results on Few-Shot In-Context Learning
Due to space limit, we only show the results of some most representative LLMs in Sec. 4.3; we show the results for the all tested LLMs in Tab. 12.
@@ C.6 More Results on Few-Shot In-Context Learning
C.7 More Results on “No Solution” Problems
Due to space limit, we only show the results of some most representative LLMs in Sec. 4.4; we show the results for the all tested LLMs in Tab. 13.
@@ C.7 More Results on “No Solution” Problems
Table 13: Results of all LLMs with the settings in Sec. $4.$ While generally adding “forced correct” prompts and 1-shot helps the performance, most models still heavily struggle.
C.8 Failure Example on Original Problem with “Forced Correct” Prompt
In Sec. 4.1, the experiment results show that with “Forced Correct” prompt, the performance of LLMs on original problems generally decrease, and such issue is particularly severe with Gemini-2.0 Pro-0205. Here we show an example of the LLM getting confused with the prompt on original problems in the box below, showing that it is non-trivial to balance correctness and user experiences with the inputs that are not so rigorous.
# Problem 请回答下面的问题。题目保证没有错误,请严格按照题目字面所写的问题回答。 【问题】你正在图书馆枯坐,一位陌生美女主动过来和你搭讪,递给你两个相同的硬币并要求和你一起玩个数学游戏。美女提议:“让我们各自亮出硬币的一面,或花或字。如果我们都是花面,那么我给你 3 元,如果我们都是字面,我给你 1 元,剩下的情况你给我 2 元就可以了。” 那么,这个游戏公平吗? (Please answer the following problems. The problems are guaranteed to be correct; please strictly follow the literal meaning of the problem. [Question] You are sitting idly in the library when a beautiful stranger approaches you and strikes up a conversation. She hands you two identical coins and proposes a math game. She says: "Let’s each reveal one side of a coin—either heads or tails. If we both show heads, I’ll give you 3 yuan. If we both show tails, I’ll give you 1 yuan. In all other cases, you just give me 2 yuan." So, is this game fair?) # Gemini-2.0 Pro-0205 Response 这个游戏是否公平,取决于 “公平” 的定义是什么。题目没有问这个游戏是否公平,也没有定义什么是 “公平”,因此无法回答这个问题。 (Whether the game is “fair” depends on the definition of fair. The problem does not ask whether the game is fair, and did not define what is “fairness”; therefore, the problem cannot be answered.)